# TUNeS: A Temporal U-Net with Self-Attention for Video-based Surgical Phase Recognition

### v3

@@ Abstract
To enable context-aware computer assistance in the operating room of the future, cognitive systems need to understand automatically which surgical phase is being performed by the medical team. The primary source of information for surgical phase recognition is typically video, which presents two challenges: extracting meaningful features from the video stream and effectively modeling temporal information in the sequence of visual features. For temporal modeling, attention mechanisms have gained popularity due to their ability to capture long-range dependencies. In this paper, we explore design choices for attention in existing temporal models for surgical phase recognition and propose a novel approach that uses attention more effectively and does not require hand-crafted constraints: TUNeS, an efficient and simple temporal model that incorporates self-attention at the core of a convolutional U-Net structure. In addition, we propose to train the feature extractor, a standard CNN, together with an LSTM on preferably long video segments, i.e., with long temporal context. In our experiments, almost all temporal models performed better on top of feature extractors that were trained with longer temporal context. On these contextualized features, TUNeS achieves state-of-the-art results on the Cholec80 and AutoLaparo datasets.
@@
{IEEEkeywords}
action segmentation, attention, Cholec80, sequence modeling, surgical phase, surgical workflow
@@ 1 Introduction
Video-based surgical phase recognition refers to automatically analyzing the video stream that is recorded in the operating room to recognize which surgical phase [1] is being performed. Robust and accurate methods for surgical phase recognition are a prerequisite for computer-assisted surgery, where context-aware assistance functions are to be provided automatically. Here, automatic computer assistance could comprise, for example, providing crucial patient-related information at the right time (intra-operatively) or preparing the surgical report (post-operatively).
@@ 1.0.3 Evaluation
We conducted a large number of experiments to demonstrate the positive effects of training the feature extractor with long temporal context (see Fig. 6). Our experiments also show that TUNeS compares favorably to many baseline methods, in terms of both recognition accuracy and computational efficiency (see Fig. 6 and Fig. 7). TUNeS, combined with a feature extractor that is trained with long temporal context, achieves state-of-the-art results on Cholec80 and AutoLaparo. Source code will be made available.
@@ 2.1 Surgical phase recognition
We denote the true phase that is happening at time $t$ as $y_{t}$, $1\leq y_{t}\leq C$, and the sequence of all ground truth labels as $\left(y_{t}\right)_{1\leq t\leq T}$ or $y$. The estimate of the phase at time $t$ is $\hat{y}_{t}\in\mathbb{R}^{C}$, where $\hat{y}_{t,p}$ is the score for the event that $y_{t}$ equals phase $p$. The video itself is a sequence of video frames $v_{t}$.
@@ 3.1 Phase recognition without attention mechanisms
Early models for video-based phase recognition consisted of a CNN for feature extraction and an LSTM network as temporal model [30, ch. 6.4]. Later, Czempiel et al. [13] proposed TeCNO for temporal modeling (Fig. 2a). TeCNO is a TCN with multiple stages [4], where subsequent stages refine predictions of previous stages. Each stage consists of several layers with dilated temporal convolutions (Fig. 1b), where the dilation factor doubles at each layer. For online recognition, causal dilated convolutions [31] are used, i.e., convolutions are computed over the current and preceding time steps only. Yi et al. [32] suggested to train multi-stage models not end-to-end, but to apply disturbances to the intermediate predictions during training. Recently, Kadkhodamohammadi et al. [33] introduced PATG, which uses a SE-ResNet as feature extractor and a graph neural network as temporal model. Zhang et al. [34] proposed a reinforcement learning approach to retrieve phase transitions for offline recognition.
@@ 4.2 Temporal U-Net with self-attention
The proposed model is a temporal U-Net with self-attention: TUNeS (Fig. 4). Previously, similar architectures were used successfully for image generation [39, 40], for image segmentation [41, 42], and for speech denoising [43].
@@ 4.2.4 Training objective
Let $\left\{\hat{y}^{(i)}:1\leq i\leq 4\right\}$ be the multi-scale phase predictions, where the length of prediction $\hat{y}^{(i)}$ is reduced by factor $\chi_{i}$ and $\chi=(1,3,9,18)$. For $2\leq i\leq 4$, we compute downsampled annotations $y^{(i)}=(y^{(i)}_{s})_{1\leq s\leq T/\chi_{i}}$, where $y^{(i)}_{s}\in\{0,1\}^{C}$ is a binary vector that summarizes the section $\{y_{t}:(s-1)\chi_{i}<t\leq s\chi_{i}\}$ in the original annotation $y$: If phase $p$ appears in the section, the $p$-th entry in $y^{(i)}_{s}$ is 1 and 0 otherwise. Since a section may cover more than one phase, $y^{(i)}_{s}$ may have more than one entry that is non-zero. Thus, the downsampled annotations are multi-label annotations.
The training objective $\mathcal{L}_{\text{total}}$ summarizes losses at all scales, see (2). We use cross-entropy (CE) loss for the full-resolution prediction $\hat{y}^{(1)}$. For the lower-resolution predictions, we use binary cross-entropy (BCE) loss to handle the multi-label annotations. To counter class imbalance, we use class weights $\gamma_{p}$, $1\leq p\leq C$, which are computed on the training data following the median frequency balancing [47] approach. In addition, we compute a smoothing loss $\mathcal{L}_{\text{smooth}}$ [4] on the full-resolution prediction, which is weighted by factor $\lambda_{\text{smooth}}=0.15$.
@@
Equations (3)–(5) define the individual loss terms:
$\mathcal{L}_{\text{CE}}(\hat{y},y)=-\frac{1}{T}\sum_{t=1}^{T}\gamma_{y_{t}}\cdot\log{S(\hat{y}_{t})_{y_{t}}}$ (3)
$\begin{split}\mathcal{L}_{\text{smooth}}(\hat{y})=\frac{1}{(T-1)C}\sum_{t=2}^{T}\sum_{p=1}^{C}\min\left(\Delta_{t,p},\tau\right)^{2},\text{ where }\\ \Delta_{t,p}=|\log{S(\hat{y}_{t})_{p}}-\mathrm{sg}(\log{S(\hat{y}_{t-1})_{p}})|\text{ and }\tau=4\end{split}$ (4)
$\begin{split}\mathcal{L}_{\text{BCE}}(\hat{y},y)=-\frac{1}{TC}\sum_{t=1}^{T}\sum_{p=1}^{C}(\,\gamma_{p}\cdot y_{t,p}\cdot\log{\left(\sigma(\hat{y}_{t,p})\right)}+\\ (1-y_{t,p})\cdot\log{\left(1-\sigma(\hat{y}_{t,p})\right)}\,)\end{split}$ (5)
Here, $\sigma$ refers to the sigmoid function, $S$ refers to the Softmax function, i.e., $S(\hat{y}_{t})_{p}=\exp{\hat{y}_{t,p}}/\sum_{c=1}^{C}\exp{\hat{y}_{t,c}}$, and $\mathrm{sg}$ refers to the $\mathrm{stopgradient}$ operator, which turns its operand into a constant with zero gradients.
@@ 5.2.1 Feature extractors
Figure 6: Surgical phase recognition results on the Cholec80 dataset (32:8:40 split). We present the feature extractor’s standalone performance as well as the performance of various temporal models, trained on top of the frozen feature extractor. Error bars indicate the standard deviation over repeated experimental runs. The asterisk* indicates training without feature sequence augmentation and smoothing loss. Numerical results can be found in the appendix (Tables 3 and 4).
5.2.2 Temporal models
For each context length $L\in\{1,8,64\}$, we trained TUNeS as well as other temporal models on the extracted feature sequences (Fig. 6). The baseline models included an RNN (2-layer GRU), an MS-TCN (TeCNO), and models that integrate attention in different ways. For comparability, we set the hidden dimension of all models to $\mathrm{dim}=64$. To account for random effects, we repeated each experiment five times on each of the five feature extractor instances, yielding 25 experimental runs in total.
To select an appropriate learning rate $\eta$ for each model, we ran a quick sweep over $\eta\in\{10^{-4},3\cdot 10^{-4},5\cdot 10^{-4},10^{-3}\}$. Usually, $\eta=5\cdot 10^{-4}$ worked well. If available, we tried to consider the training strategy of the original code base. Full training details are listed in appendix .1.
@@ 5.2.2 Temporal models
5.2.3 Performance measurements
@@
5.2.4 Comparison to the state of the art
We compare the results of TUNeS, trained on features with maximum context $L=64$, with previously reported results on the 40:40 split (Tables 1 and 2). Here, the feature extractor with $L=64$ and TUNeS were re-trained on all of the first 40 videos in Cholec80. Due to the lack of a validation set, we selected the model after the last epoch for testing. To improve convergence, we continuously reduced the learning rate (lr) during training, following a cosine annealing schedule [52]. For comparisons on the 40:20:20 split, we evaluated the models trained for the 40:40 split on the last 20 videos in Cholec80.
Discussion
When evaluating the baselines (Fig. 6), we found that the 2-layer GRU and OperA achieved subpar results. For the GRU, we suspect that regularization with the smoothing loss and the proposed feature sequence augmentation is not as effective for a recurrent model. For OperA, we presume that the model is suffering from a lack of local inductive bias and struggles to filter information from noisy features. Yet, OperA improves considerably on contextualized features ($L=64$), which might be more meaningful and thus easier to analyze with attention. Please note that we used a custom implementation of OperA because official code has not been released yet.
The best performing baselines were SAHC, TeCNO, TUT, and ASFormer. TUNeS outperformed the baselines in terms of accuracy and scored better or similarly on Macro Jaccard (Fig. 6, Tables 3 and 4). The results achieved by TUNeS, trained on top of the feature extractor with $L=64$, also compare favorably to previous results from the literature (Tables 1 and 2). Moreover, TUNeS is computationally efficient and scales well on long sequences of up to two hours, regarding both latency and memory consumption (Fig. 7).
@@ 5.3 Ablation studies
For further insights into TUNeS, additional experiments were conducted on the 32:8:40 split.
@@ 5.3.2 Impact of alternate attention masking
We trained and tested models for offline recognition either with attention masks, which alternately prevent access to future and to past information, or without any attention masks. Here, we studied three temporal models with global attention: OperA, SAHC, and TUNeS. For SAHC, which uses individual Transformer blocks instead of stacks of multiple Transformer blocks (Fig. 2e), we duplicated each Transformer block to implement alternate attention masking.
@@ 5.3.3 Quality of predicted phase segments
The frame-wise annotation $y$ induces a partitioning into $S$ phase segments $(Y_{s})_{1\leq s\leq S}$, where each segment summarizes all consecutive frames of a phase. Likewise, the frame-wise prediction $\hat{y}$ induces the estimated segmentation $\tilde{Y}$. Due to the incorrect classification of individual frames, it is common for $\tilde{Y}$ to divide a video into many fragmented segments. To quantify this, Lea et al. [53] proposed the edit score, which is the normalized Levenshtein distance between $Y$ and $\tilde{Y}$. This metric penalizes out-of-order predictions and over-segmentation errors.
Fig. 10 shows the mean video-wise edit score for phase segments predicted in offline mode by TUNeS and other temporal models. TUNeS obtains higher edit scores than GRU, OperA, and SAHC. Yet, the 2-stage models achieve better edit scores, especially TUT. We assume that the multi-stage setup – in addition to the repeated down- and upsampling in TUT – helps to reduce noise and segment fragmentation. Consequently, we also trained a 2-stage variant of TUNeS, see (6). Here, we optimized the sum of the losses at both stages.
@@ 5.3.4 Additional ablations
Discussion
As shown in Fig. 11, most modifications impaired the results, but in many cases the decline was not severe. (i) Omitting the convolutional positional encoding did not lead to considerably worse results. Probably, the convolutions in the encoder provide sufficient local inductive bias. Also, sinusoidal positional encodings did not improve results. In online mode, they worked as well as using no positional encodings. In offline mode, they caused training instability and overall inferior results. Presumably, this kind of rigid and absolute positional encoding is not suitable for modeling past and future information in sequences of widely varying length. (ii) Foregoing dilations in the decoder convolutions slightly affected the results for online recognition. The offline model seems to be robust in this regard, possibly due to the acausal operations, which have a symmetric receptive field. (iii) Regarding different regularization techniques, we found that deactivating dropout had only a minimal effect on online and offline recognition. Dropping token masks worsened the results slightly. Omitting feature sequence augmentation had a noticeable negative effect. In particular, Macro Jaccard dropped by 1.8 % for online recognition and by 1.1 % for offline recognition. (iv) Interestingly, training TUNeS on LSTM features instead of CNN features impaired the results considerably. Possibly, some of the detail in the CNN features is lost during additional processing with the LSTM. Further, the LSTM features might fit the training data too well, making it difficult for the temporal model to learn generalizable patterns based on that. (v) The impact of attention was the greatest: Without attention to model global temporal relationships, Macro Jaccard dropped by 2.4 % in online mode and by 4.1 % in offline mode. Without attention, TUNeS performs similarly to TeCNO, which might indicate that multi-scale modeling alone cannot explain the performance of TUNeS. Rather, the effective integration of Transformer blocks is a crucial factor. (vi) When scaling the number of Transformer blocks (Fig. 12), we found that using more Transformer blocks was helpful for offline recognition but not for online recognition. Generally, it might be difficult to further improve on the online task due to the inherent uncertainty and ambiguity of this problem. Here, even global information is incomplete because it is only information from the past.
@@ 5.4 Evaluation on the AutoLaparo dataset
We additionally evaluated TUNeS on AutoLaparo [54], which consists of 21 video recordings of laparoscopic hysterectomies. This dataset provides manual annotations of $C=7$ surgical phases at a resolution of 1 fps. Following the proposed data split, we used the first 10 videos for training, the following 4 videos for validation, and the last 7 videos for testing.

### v4

@@ Abstract
Objective: To enable context-aware computer assistance in the operating room of the future, cognitive systems need to understand automatically which surgical phase is being performed by the medical team. The primary source of information for surgical phase recognition is typically video, which presents two challenges: extracting meaningful features from the video stream and effectively modeling temporal information in the sequence of visual features. Methods: For temporal modeling, attention mechanisms have gained popularity due to their ability to capture long-range dependencies. In this paper, we explore design choices for attention in existing temporal models for surgical phase recognition and propose a novel approach that uses attention more effectively and does not require hand-crafted constraints: TUNeS, an efficient and simple temporal model that incorporates self-attention at the core of a convolutional U-Net structure. In addition, we propose to train the feature extractor, a standard CNN, together with an LSTM on preferably long video segments, i.e., with long temporal context. Results: In our experiments, almost all temporal models performed better on top of feature extractors that were trained with longer temporal context. On these contextualized features, TUNeS achieves state-of-the-art results on the Cholec80 dataset. Conclusion: This study offers new insights on how to use attention mechanisms to build accurate and efficient temporal models for surgical phase recognition. Significance: Implementing automatic surgical phase recognition is essential to automate the analysis and optimization of surgical workflows and to enable context-aware computer assistance during surgery, thus ultimately improving patient care.
@@ 1 Introduction
Video-based surgical phase recognition refers to automatically analyzing the video stream that is recorded in the operating room to recognize which surgical phase [1] is being performed. Robust and accurate methods for surgical phase recognition are a prerequisite for computer-assisted surgery, where context-aware assistance functions are to be provided automatically [2]. Here, automatic computer assistance could comprise, for example, providing crucial patient-related information at the right time (intra-operatively) or preparing the surgical report (post-operatively).
@@ 1.0.3 Evaluation
We conducted a large number of experiments to demonstrate the positive effects of training the feature extractor with long temporal context (see Fig. 6). Our experiments also show that TUNeS compares favorably to many baseline methods, in terms of both recognition accuracy and computational efficiency (see Fig. 6 and Fig. 7). TUNeS, combined with a feature extractor that is trained with long temporal context, achieves state-of-the-art results on Cholec80 and AutoLaparo. The source code is available at https://gitlab.com/nct_tso_public/tunes.
@@ 2.1 Surgical phase recognition
We denote the true phase that is happening at time $t$ as $y_{t}$, $1\leq y_{t}\leq C$, and the sequence of all ground truth labels as $\left(y_{t}\right)_{1\leq t\leq T}$ or $y$. The estimate of the phase at time $t$ is $\hat{y}_{t}\in\mathbb{R}^{C}$, where $\hat{y}_{t,p}$ is the score (i.e., the logarithm of the unnormalized probability) for the event that $y_{t}$ equals phase $p$. The video itself is a sequence of video frames $v_{t}$.
@@ 3.1 Phase recognition without attention mechanisms
Early models for video-based phase recognition consisted of a CNN for feature extraction and an LSTM network as temporal model [31, chapter 6.4]. Later, Czempiel et al. [14] proposed TeCNO for temporal modeling (Fig. 2a). TeCNO is a TCN with multiple stages [5], where subsequent stages refine predictions of previous stages. Each stage consists of several layers with dilated temporal convolutions (Fig. 1b), where the dilation factor doubles at each layer. For online recognition, causal dilated convolutions [32] are used, i.e., convolutions are computed over the current and preceding time steps only. Yi et al. [33] suggested to train multi-stage models not end-to-end, but to apply disturbances to the intermediate predictions during training. Recently, Kadkhodamohammadi et al. [34] introduced PATG, which uses a SE-ResNet as feature extractor and a graph neural network as temporal model. Zhang et al. [35] proposed a reinforcement learning approach to retrieve phase transitions for offline recognition.
@@ 4.2 Temporal U-Net with self-attention
The proposed model is a temporal U-Net with self-attention: TUNeS (Fig. 4). Previously, similar architectures were used successfully for image generation [40, 41], image segmentation [42, 43], or speech denoising [44].
@@ 4.2.4 Training objective
Let $\left\{\hat{y}^{(i)}:1\leq i\leq 4\right\}$ be the multi-scale phase predictions, where the length of prediction $\hat{y}^{(i)}$ is reduced by factor $\chi_{i}$ and $\chi=(1,3,9,18)$. For $2\leq i\leq 4$, we compute downsampled annotations $y^{(i)}=(y^{(i)}_{s})_{1\leq s\leq T/\chi_{i}}$, where $y^{(i)}_{s}\in\{0,1\}^{C}$ is a binary vector that summarizes the section $\{y_{t}:(s-1)\chi_{i}<t\leq s\chi_{i}\}$ of $\chi_{i}$ frames in the full-resolution annotation: If phase $p$ appears in the section, the $p$-th entry in $y^{(i)}_{s}$ is 1 and 0 otherwise. Since a section may cover more than one phase, $y^{(i)}_{s}$ may have more than one entry that is non-zero. Thus, the downsampled annotations are multi-label annotations.
The training objective $\mathcal{L}_{\text{total}}$ summarizes losses at all scales, see (2). We use cross-entropy (CE) loss for the full-resolution prediction $\hat{y}^{(1)}$. For the lower-resolution predictions, we use binary cross-entropy (BCE) loss to handle the multi-label annotations. To counter class imbalance, we use class weights $\gamma$, which are computed on the training data following the median frequency balancing [48] approach. In addition, we compute a smoothing loss $\mathcal{L}_{\text{smooth}}$ [5] on the full-resolution prediction, which is weighted by factor $\lambda_{\text{smooth}}=0.15$.
@@
The individual loss terms are defined in the supplementary.
@@ 5.2.1 Feature extractors
5.2.2 Temporal models
For each context length $L\in\{1,8,64\}$, we trained TUNeS as well as other temporal models on the extracted feature sequences (Fig. 6). The baseline models included an RNN (2-layer GRU), an MS-TCN (TeCNO), and models that integrate attention in different ways. For comparability, we set the hidden dimension of all models to $\mathrm{dim}=64$. To account for randomness, we repeated each experiment five times on each of the five feature extractor instances, yielding 25 experimental runs in total.
To select an appropriate learning rate $\eta$ for each model, we ran a quick sweep over $\eta\in\{10^{-4},3\cdot 10^{-4},5\cdot 10^{-4},10^{-3}\}$. Usually, $\eta=5\cdot 10^{-4}$ worked well. If available, we tried to consider the training strategy of the original code base. Full training details are listed in the supplementary material.
@@ 5.2.2 Temporal models
Figure 6: Surgical phase recognition results on the Cholec80 dataset (32:8:40 split). We present the feature extractor’s standalone performance as well as the performance of various temporal models, trained on top of the frozen feature extractor. Error bars indicate the standard deviation over repeated experimental runs. The asterisk* indicates training without feature sequence augmentation and smoothing loss. Numerical results can be found in the supplementary material.
5.2.3 Performance measurements
5.2.4 Comparison to the state of the art
We compare the results of TUNeS, trained on features with maximum context $L=64$, with previously reported results on the 40:40 split (Tables 1 and 2). Here, the feature extractor with $L=64$ and TUNeS were re-trained on all of the first 40 videos in Cholec80. Due to the lack of a validation set, we selected the model after the last epoch for testing. To improve convergence, we continuously reduced the learning rate during training, following a cosine annealing schedule [53]. For comparisons on the 40:20:20 split, we evaluated the models trained for the 40:40 split on the last 20 videos in Cholec80.
@@ 5.2.4 Comparison to the state of the art
Discussion
When evaluating the baselines (Fig. 6), we found that the 2-layer GRU and OperA achieved subpar results. For the GRU, we suspect that regularization with the smoothing loss and the proposed feature sequence augmentation is not as effective for a recurrent model. For OperA, we presume that the model is suffering from a lack of local inductive bias and struggles to filter information from noisy features. Yet, OperA improves considerably on contextualized features ($L=64$), which might be more meaningful and thus easier to analyze with attention. Please note that we used a custom implementation of OperA because official code has not been released.
The best performing baselines were SAHC, TeCNO, TUT, and ASFormer. TUNeS outperformed the baselines in terms of accuracy and scored better or similarly on Macro Jaccard (Fig. 6). The results achieved by TUNeS, trained on top of the feature extractor with $L=64$, also compare favorably to previous results from the literature (Tables 1 and 2). Moreover, TUNeS is computationally efficient and scales well on long sequences of up to two hours, regarding both latency and memory consumption (Fig. 7).
@@ 5.3 Ablation studies
For further insights into TUNeS, additional experiments were conducted on the 32:8:40 split. In each ablation, one specific aspect was varied as described below.
@@ 5.3.2 Impact of alternate attention masking
We trained and tested models for offline recognition either with attention masks, which alternately prevent access to future and to past information, or without any attention masks. Here, we studied three temporal models with global attention: OperA, SAHC, and TUNeS. For SAHC, which uses individual Transformer blocks instead of stacks of multiple Transformer blocks (Fig. 2e), we duplicated each Transformer block to implement alternate attention masking. Then, the first block queries past information and the second block queries future information.
@@ 5.3.3 Quality of predicted phase segments
The frame-wise annotation $y$ induces a partitioning $Y$ into phase segments, where each segment summarizes all consecutive frames of a phase. Likewise, the frame-wise prediction $\hat{y}$ induces the estimated segmentation $\tilde{Y}$. Due to the incorrect classification of individual frames, it is common for $\tilde{Y}$ to divide a video into many fragmented segments. To quantify this, Lea et al. [54] proposed the edit score, which is the normalized Levenshtein distance between $Y$ and $\tilde{Y}$. This metric penalizes out-of-order predictions and over-segmentation errors.
Fig. 10 shows the mean video-wise edit score for phase segments predicted in offline mode by TUNeS and other temporal models. TUNeS obtains higher edit scores than GRU, OperA, and SAHC. Yet, the 2-stage models yield better segmentations, especially TUT. We assume that the multi-stage setup – in addition to the repeated down- and upsampling in TUT – helps to reduce noise and segment fragmentation. Consequently, we also trained a 2-stage variant of TUNeS, see (3). Here, we optimized the sum of the losses at both stages.
@@ 5.3.4 Additional ablations
Discussion
As shown in Fig. 11, most modifications impaired the results, but in many cases the decline was not severe.
(i) Omitting the convolutional positional encoding did not lead to considerably worse results. Probably, the convolutions in the encoder provide sufficient local inductive bias. Also, sinusoidal positional encodings did not improve results. In online mode, they worked as well as using no positional encodings. In offline mode, they caused training instability and overall inferior results. Presumably, this kind of rigid and absolute positional encoding is not suitable for modeling past and future information in sequences of widely varying length. (ii) Foregoing dilations in the decoder convolutions slightly affected the results for online recognition. The offline model seems to be robust in this regard, possibly due to the acausal operations, which have a symmetric receptive field. (iii) Regarding different regularization techniques, we found that deactivating dropout had only a minimal effect on online and offline recognition. Dropping token masks worsened the results slightly. Omitting feature sequence augmentation had a noticeable negative effect. In particular, Macro Jaccard dropped by 1.8 % for online recognition and by 1.1 % for offline recognition. (iv) Interestingly, training TUNeS on LSTM features instead of CNN features impaired the results considerably. Possibly, some of the detail in the CNN features is lost during additional processing with the LSTM. Further, the LSTM features might fit the training data too well, making it difficult for the temporal model to learn generalizable patterns based on that. (v) The impact of attention was the greatest: Without attention to model global temporal relationships, Macro Jaccard dropped by 2.4 % in online mode and by 4.1 % in offline mode. Without attention, TUNeS performs similarly to TeCNO, which might indicate that multi-scale modeling alone cannot explain the performance of TUNeS. Rather, the effective integration of Transformer blocks is a crucial factor. (vi) When scaling the number of Transformer blocks (Fig. 12), we found that using more Transformer blocks was helpful for offline recognition but not for online recognition. Generally, it might be difficult to further improve on the online task due to the inherent uncertainty and ambiguity of this problem. Here, even global information is incomplete because it is only information from the past.
@@ 5.4 Evaluation on the AutoLaparo dataset
We additionally evaluated TUNeS on AutoLaparo [^2] https://autolaparo.github.io [55], which consists of 21 video recordings of laparoscopic hysterectomies. This dataset provides manual annotations of $C=7$ surgical phases at a resolution of 1 fps. Following the proposed data split, we used the first 10 videos for training, the following 4 videos for validation, and the last 7 videos for testing.